cube with edge length a is h filled with water of density p. Find the force exerted on a side of the cube by the water. (The edges of the cube are either hor- izontal or vertical.)
Answer #1

Here , for the cube ,

pressure at the top surface of water, P1 = 0 Pa

Pressure at the bottom of water , P2 = p * g * a/2

Now, for the average pressure , P = (P1 + P2)/2

P = (0 + pga/2)/2

P = pga/4

force exerted on the side = average pressure * area

force exerted on the side = pga/4 * a^2/2

force exerted on the side = pga^3/8

force exerted on the side is p *g*a3/8

where p is the density of water